# Three-dimensional shape and connectivity of physical networks

### v1

@@ Abstract
Data describing the three-dimensional structure of physical networks is increasingly available, leading to a surge of interest in network science to explore the relationship between the shape and connectivity of physical networks. We contribute to this effort by standardizing and analyzing 15 data sets from different domains. Treating junction points as nodes and connections between them as links, we divide the networks into three categories: lattice-like networks, trees, and linked trees. We find that the degree distribution of physical networks is bounded, with most nodes having degree one or three. Characterizing the physical properties of links, we show that links have an elongated shape and tend to follow a nearly straight trajectory, while a small fraction of links follow a winding path. These typical node and link properties must be reflected by physical network models. We also measure how confined a link is in space by comparing its trajectory to a randomized null-model, showing that links that are central in the abstract network tend to be physically confined by their neighbors. The fact that the shape and connectivity of the physical networks are intertwined highlights that their three-dimensional layout must be taken into account to understand the evolution and function of physical networks.
@@ I Introduction
Recently available detailed maps of physical networks provide an opportunity to systematically investigate the relationship between their shape and their network properties. For example, open-data large-scale experimental projects provide a neuron-level mapping of biological neural networks [34, 26], high-throughput MRI measurements map out detailed vascular networks [10], or mycelia mapping projects explore local morphological and mechanical properties of fungi by representing them as networks of filaments [11]. The nodes and links of these physical networks have two distinguishing features: (i) they are characterized by a complex three-dimensional shape and (ii) they physically interact with each other, for example, nodes and links obey volume exclusion, i.e., they cannot overlap. To understand how physicality affects network evolution and function, we must extend the toolset of network science to take into account these features.
@@
Recent work investigated artificial spatial embeddings of complex networks that obey volume exclusion [6], the entanglement of physical links [15], models of physical network growth [20, 19], and the effect of physical shape on the dynamics on networks [19]. However, systematic exploration of the three-dimensional shape, the network properties and the relationship between them in real networks is still lacking. Such exploration is hampered by the lack of standardized representation. First, there is the technical difficulty that experimental maps of physical networks, like neural or molecular networks, are collected, processed, and analyzed with domain-specific methodology. Therefore, any investigation of physical networks must be preceded by the time-consuming and computationally burdensome task of data pre-processing. Second, even seemingly simple questions like what is a node and a link in a physical network carry a level of ambiguity: a physical network is a continuous object in space; to represent it as a network, we must discretize it into nodes and links. The definition of nodes and links in turn affects, for example, what properties of the network we can study or the right choice of null models.
@@
Here, we compile and standardize 15 data sets from various domains. Each of these physical networks is composed of tube-like objects bound together at junction points; motivating us to treat the junction points as nodes and the tubes connecting them as physical links. We characterize both physical shape and the abstract network structure and the correlations between them. For this, we calculate standard descriptors such as the degree distribution of the abstract network or the fractal dimension of the layout. We also introduce a measure of link confinement to understand the role of volume exclusion, which compares the physical links to a null model that randomizes link routing. The remainder of the paper is organized as follows: In the next section, we describe the data sets we collected and their standardization. In. Secs. III-V, we analyze the data sets’ abstract network properties, their physical shape, and the emergent correlations between network and shape. Finally, Sec. VI provides a brief discussion.
@@
Figure 1: Physical networks. (a) A physical network composed of three tube-like links bound together at a single junction point. (b) The skeleton representation of (a) approximates the original structure as a collection of vertices (black points) and edges (colored segments). (c) The combinatorial or abstract network of (a) captures the connectivity of the system without the physical structure: nodes represent junction points and terminal points, with a link between two the nodes that they are directly connected by a physical link. (d) The skeleton representation allows us to approximate the original volume of real physical networks – e.g., vascular network (top) and mitochondrial network (bottom). (e) We compiled a set of 15 physical networks from various domains. The size of the networks varies greatly: the number of skeleton segments $N_{\textnormal{seg}}$ capturing the shape of each networks spans 4 orders of magnitude.
@@ II Data
Our goal is to systematically study the three-dimensional layout of physical networks and to understand the relationship between their physical properties and their network structure. For this, we collected 15 data sets from various domains, including individual neurons [23] [3] [12] [30], biological neural networks [25], plant roots [17], vascular networks [10], a mitochondrial network [31], and the imprint of an anthill (See SI Sec. I). Before any analysis, however, we must uniformly represent and standardize these data sets. In the following sections, we propose the use of a labeled skeleton representation, which efficiently captures both the physical shape and abstract network of physical networks.
@@ II.1 Skeleton representation
Physical networks are typically composed of tube-like objects bound together at junction points (Fig. 1a), making them suitable to be approximated by a series of straight segments and radii of the network at the endpoints of the segments (Fig. 1b). The process of creating these segments from a raw data format is called skeletonization [22], often employed in the fields of computer graphics [29] and neuroscience [27, 2]. Formally, a skeleton representation ${\mathcal{S}}$ of a physical network is a graph whose set of vertices ${\mathcal{V}}$ correspond to points in space and set of edges ${\mathcal{E}}$ correspond to segments connecting point pairs. Therefore each vertex $i\in{\mathcal{V}}$ has a set of coordinates ${\bf r}_{i}=(x_{i},y_{i},z_{i})$ and a radius $\rho_{i}$ associated to it. Figure 1b shows the skeleton of Fig. 1a, the black circles represent the skeleton vertices.
The radius $\rho_{i}$ captures the thickness of the physical network at each vertex $i$. Therefore, we can approximate the original occupied volume of the network as a union of truncated cones: for each edge $(i,j)$ in ${\mathcal{S}}$ we add a truncated cone with axis $({\bf r}_{i},{\bf r}_{j})$ and parallel faces with radii $\rho(i)$ and $\rho(j)$ which has volume:
@@
    Merging segments: If two consecutive segments $({\bf r}_{i},{\bf r}_{j})$ and $({\bf r}_{j},{\bf r}_{k})$ are approximately parallel to each other, we merge them into a single segment $({\bf r}_{i},{\bf r}_{k})$ (See SI Sec. II B).
@@
    Skeleton healing: Due to noisy data, a skeleton may be disconnected even when it represents a single continuous object in reality. For example, skeleton of a neuron may appear to have multiple components. To remedy this, we add a segment to connect the two nearest skeleton vertices from two disconnected components. We repeat this step until the skeleton becomes connected.
@@ II.2 Network structure
A physical network is a continuous object embedded in Euclidean space, to characterize this object as a network we must separate it into discrete nodes and links. For this, first note that all 15 data sets that we collected can be seen as a collection of tube-like objects bound together at junction points. Motivated by this observation, we define the junction and terminal points of the tubes as physical nodes and the non-branching tubes pairwise connecting these terminal and junction points as physical links. A motivation for this definition is that cutting a physical link (i.e., a tube) at any point along its length causes the same disruption to the connectivity of the network.
@@ III Abstract network properties
The nodes in our physical networks are terminal and junction points, meaning that by construction nodes cannot have degree $k=2$, only degree $k=1$ or $k>2$. Figure 2a shows the degree distribution $p(k)$ of all 15 networks, our main observation is that most nodes have degree $k=1$ or $k=3$, and nodes with a larger degree are exceedingly rare, i.e., $p(1)+p(3)\approx 1$ This means that junctions tend to be bifurcation points along the tube-like physical links making up the network. This is in line with previous empirical observations and theoretical predictions for neurons [18, 7] and transport networks [13, 8]. The observed narrow degree distribution is in contrast with degree-heterogeneous networks typically in the focus of network science and should be accounted for by mathematical models of physical networks [20].
@@
To explore the local loop structure of the networks, we calculate the abundance of observed 4-node motifs. Namely, we focus on two motifs: the star motif and the 4-cycle. To quantify their abundance, we calculate their Z-score compared to their degree preserving randomized counterparts:
@@
where $n_{\textrm{s/c}}$ is the number of occurrences of the star and 4-cycle motifs in the original networks, and the expected value $\left\langle n_{\textrm{s/c}}\right\rangle$ and standard deviation $\sigma_{\textrm{s/c}}$ is estimated by creating 200 independent randomizations[16], while keeping the degree sequence fixed. Figure 2b shows the scatter plot of $z_{\textrm{s}}$ and $z_{\textrm{c}}$ for the 15 networks. As expected, in tree networks containing no cycles the star motif is slightly over-represented $z_{\textrm{s}}>0$, and the cycle motif is slightly underrepresented $z_{\textrm{s}}<0$, while for networks containing cycles, we find the opposite. The highest abundance of 4-cycles is observed for the networks representing different brain regions of the fruit fly brain.
@@
Figure 2: Abstract network properties (a) The abstract networks are composed of terminal and bifurcation points; therefore, their degree distribution is mostly concentrated on $k=1$ and $k=3$. More specifically, lattice-like networks, such as vascular and mitochondrial networks, are mostly made up of branching nodes (degree $k=3$). For the rest of the network, nodes with $k=1$ and $k=3$ are approximately evenly split, as expected for tree networks. (b) We calculate the $z$-scores of four-node star and cycle motifs of the original networks compared to random networks with the same degree distribution. Markers are the average of $n_{\textnormal{t}}=200$ independent randomizations, error bars representing the standard error of the mean are smaller than the marker size. Both axes are symlog axes, i.e., linear from -1 to 1, log otherwise for both positive and negative values. The star motif is slightly over and the cycle motif is slightly underrepresented in tree networks (green markers), as expected. In contrast. In contrast, stars are under- and cycles are over-represented in both lattices (blue markers) and linked trees (red markers). The high $z$-scores observed for the 4-cycle indicate a lattice-like structure.
@@ IV Physical properties
In the previous section, we measured properties of the abstract network capturing connectivity in the 15 physical networks. We continue our investigation focusing on the physical properties of the system: we characterize the three-dimensional shape of the system without considering the abstract network.
@@ IV.1 Space filling and fractal dimension
A fundamental property of a physical network is its density or space-filling, i.e., the amount of volume it occupies from the available space. We expect that networks that are tightly packed in space are strongly affected by physicality [6, 21, 19]. Here, we measure a local space-filling of the 15 physical networks by dividing the bounding box of each physical network into cubes with side length $l_{b}$ and calculating the space-filling within each cube $i$ as
$\phi(i)=\frac{V_{\textrm{occ}}(i)}{l_{b}^{3}},$ (3)
@@
where $V_{\textrm{occ}}(i)$ is the volume of the intersection of the networks and cube $i$. The distribution of $\phi(i)$ depends on the choice of $l_{b}$; therefore, to ensure consistency, we set $l_{b}$ separately for each dataset such that the cubes are most likely to contain one or two links (For details, check SI. Sec LABEL:physical_density_SI). Figure 3 shows that the distribution of $\phi$ for all 15 physical networks is not peaked around a single value, but the physical networks fill out the space unevenly with both densely packed and sparse regions. Therefore we expect that physicality will also affect the network structure unevenly: volume exclusion can limit the number and shape of links in dense regions.
@@
To further characterize the shape of the networks, we calculate their box-counting fractal dimension $D_{\textnormal{f}}$, which compactly describes the space-filling of a physical object on multiple scales [28, 9, 32], and is widely used to characterize the shape of complex biological systems [32, 4]. Possible values of $D_{\textnormal{f}}$ for connected networks range between 1 and the embedding dimension $D=3$, Fig. 3b shows $D_{\textnormal{f}}$ for the 15 networks, we find that both lattices and linked trees have $D_{\textnormal{f}}\gtrapprox 2$, while trees are typically characterized by $D_{\textnormal{f}}\lessapprox 2$, except for the anthill imprint. The fractal scaling spans at least two orders of magnitude of length scales (see SI Sec.LABEL:fractal_dimension_SI), again pointing towards regions of high and low physical density at different resolution levels.
@@
Figure 3: Space-filling and fractal dimension. (a) The distribution of local space-filling $\phi$ shows the coexistence of mostly sparse and fewer dense regions. (b) The fractal dimension of physical networks $D_{\textrm{f}}$ ranges between $1$ and the embedding dimension $D=3$. Linked trees (red) and lattices (blue) have fractal dimension values between $D_{\textrm{f}}\approx 2.0$ and $D_{\textrm{f}}\approx 2.3$, while trees (green) have more variation, ranging from $D_{\textrm{f}}\approx 1.5$ to $D_{\textrm{f}}\approx 2.0$ (except for from the anthill imprint). We estimate $D_{\textrm{f}}$ using the box-counting method, the error bars indicate the standard deviation of the local scaling estimates (see SI Sec.LABEL:fractal_dimension_SI).

### v2

@@ Abstract
Data describing the three-dimensional structure of physical networks is increasingly available, leading to a surge of interest in network science to explore the relationship between the shape and connectivity of physical networks. We contribute to this effort by standardizing and analyzing 15 data sets from different domains. Each network is made of tube-like objects bound together at junction points, which we treat as nodes, with the connections between them considered as links. We divide these networks into three categories: lattice-like networks, trees, and linked trees. The degree distribution of these physical networks is bounded, with most nodes having degrees one or three. Characterizing the physical properties of links, we show that links have an elongated shape and tend to follow a nearly straight trajectory, while a small fraction of links follow a winding path. These typical node and link properties must be reflected by physical network models. We also measure how confined a link is in space by comparing its trajectory to a randomized null model, showing that links that are central in the abstract network tend to be physically confined by their neighbors. The fact that the shape and connectivity of the physical networks are intertwined highlights that their three-dimensional layout must be taken into account to understand the evolution and function of physical networks.
@@ I Introduction
Recently available detailed maps of physical networks provide an opportunity to systematically investigate the relationship between their shape and their network properties. For example, open-data large-scale experimental projects provide a neuron-level mapping of biological neural networks [1, 2], high-throughput MRI measurements map out detailed vascular networks [3], or mycelia mapping projects explore local morphological and mechanical properties of fungi by representing them as networks of filaments [4]. The nodes and links of these physical networks have two distinguishing features: (i) they are characterized by a complex three-dimensional shape and (ii) they physically interact with each other, for example, nodes and links obey volume exclusion, i.e., they cannot overlap. To understand how physicality affects network evolution and function, we must extend the toolset of network science to take these features into account.
@@
Recent work investigated artificial spatial embeddings of complex networks that obey volume exclusion [5], the entanglement of physical links [6], models of physical network growth [7, 8], and the effect of physical shape on the dynamics on networks [8]. However, systematic exploration of the three-dimensional shape, the network properties, and the relationship between them in real networks is still lacking. Such exploration is hampered by the lack of standardized representation. First, there is the technical difficulty that experimental maps of physical networks, like neural or molecular networks, are collected, processed, and analyzed with domain-specific methodology. Therefore, any investigation of physical networks must be preceded by the time-consuming and computationally burdensome task of data pre-processing. Second, even seemingly simple questions like what is a node and a link in a physical network carry a level of ambiguity: a physical network is a continuous object in space; to represent it as a network, we must discretize it into nodes and links. The definition of nodes and links in turn affects, for example, what properties of the network we can study or the right choice of null models.
@@
Here, we compile and standardize 15 data sets from various domains. Each of these physical networks is composed of tube-like objects bound together at junction points; motivating us to treat the junction points as nodes and the tubes connecting them as physical links. We characterize both the physical shape and the abstract network structure and the correlations between them. For this, we calculate standard descriptors such as the degree distribution of the abstract network or the fractal dimension of the layout. We also introduce a measure of link confinement to understand the role of volume exclusion, which compares the physical links to a null model that randomizes link trajectories. The remainder of the paper is organized as follows: In the next section, we describe the data sets we collected and their standardization. In sections III-V, we analyze the data sets’ abstract network properties, their physical shape, and the emergent correlations between network and shape. Finally, Sec. VI provides a brief discussion.
@@
Figure 1: Physical networks. (a) A physical network composed of three tube-like links bound together at a single junction point. (b) The skeleton representation of (a) approximates the original structure as a collection of vertices (black points) and edges (colored segments). (c) The combinatorial or abstract network of (a) captures the connectivity of the system without the physical structure: nodes represent junction points and terminal points, with a link between two nodes that are directly connected by a physical link. (d) The skeleton representation allows us to approximate the original volume of real physical networks – e.g., vascular network (top) and mitochondrial network (bottom). (e) We compiled a set of 15 physical networks from various domains. The size of the networks varies greatly: the number of skeleton segments $N_{\textnormal{seg}}$ capturing the shape of each network spans 4 orders of magnitude.
@@ II Data
Our goal is to systematically study the three-dimensional layout of physical networks and to understand the relationship between their physical properties and their network structure. For this, we collected 15 data sets from various domains, including individual neurons [9, 10, 11, 12], biological neural networks [13], plant roots [14], vascular networks [3], a mitochondrial network [15], and the imprint of an anthill (see SI S1). Before any analysis, however, we must uniformly represent and standardize these data sets. In the following sections, we propose the use of a labeled skeleton representation, which efficiently captures both the physical shape and abstract network of physical networks.
@@ II.1 Skeleton representation
Physical networks are typically composed of tube-like objects bound together at junction points (Fig. 1a), making them suitable to be approximated by a series of straight segments and radii of the network at the endpoints of the segments (Fig. 1b). The process of creating these segments from a raw data format is called skeletonization [16], often employed in the fields of computer graphics [17] and neuroscience [18, 19]. Formally, a skeleton representation ${\mathcal{S}}$ of a physical network is a graph whose set of vertices ${\mathcal{V}}$ correspond to points in space and set of edges ${\mathcal{E}}$ correspond to segments connecting point pairs. Therefore each vertex $i\in{\mathcal{V}}$ has a set of coordinates ${\bf r}_{i}=(x_{i},y_{i},z_{i})$ and a radius $\rho_{i}$ associated to it. Figure 1b shows the skeleton of Fig. 1a.
The radius $\rho_{i}$ captures the thickness of the physical network at each vertex $i$. Therefore, we can approximate the original occupied volume of a segment connecting vertex $i$ with radius $\rho_{i}$ and vertex $j$ with $\rho_{j}$ as a truncated cone, and the full volume of the network is approximated by the sum of these truncated cones. More specifically, for each edge $(i,j)$ in ${\mathcal{S}}$ we add a truncated cone with axis $({\bf r}_{i},{\bf r}_{j})$ and parallel faces with radii $\rho_{i}$ and $\rho_{j}$ which has volume:
@@
    Merging segments: If two consecutive segments $({\bf r}_{i},{\bf r}_{j})$ and $({\bf r}_{j},{\bf r}_{k})$ are appear parallel to each other in the data set, we merge them into a single segment $({\bf r}_{i},{\bf r}_{k})$ (see SI S2.2).
@@
    Skeleton healing: Due to noisy data, a skeleton may be disconnected even when it represents a single continuous object in reality. For example, a skeleton of a neuron may appear to have multiple components. To remedy this, we add a segment to connect the two nearest skeleton vertices from two disconnected components. We repeat this step until the skeleton becomes connected.
@@ II.2 Network structure
A physical network is a continuous object embedded in Euclidean space, to characterize this object as a network we must separate it into discrete nodes and links. For this, first note that all 15 data sets that we collected can be seen as a collection of tube-like objects bound together at junction points. Hence, we define the junction and terminal points of the tubes as physical nodes and the non-branching tubes pairwise connecting these terminal and junction points as physical links. A motivation for this definition is that cutting a physical link (i.e., a tube) at any point along its length causes the same disruption to the connectivity of the network.
@@ III Abstract network properties
The nodes in our physical networks are terminal and junction points, meaning that by construction nodes cannot have degree $k=2$, only degree $k=1$ or $k>2$. Figure 2a shows the degree distribution $p(k)$ of all 15 networks, and our main observation is that most nodes have degree $k=1$ or $k=3$, and nodes with a larger degree are exceedingly rare, i.e., $p(1)+p(3)\approx 1$. This means that junctions tend to be bifurcation points along the tube-like physical links making up the network. This is in line with previous empirical observations and theoretical predictions for neurons [21, 22] and transport networks [23, 24]. The observed narrow degree distribution is in contrast with degree-heterogeneous networks typically in the focus of network science and should be accounted for by mathematical models of physical networks [7].
@@
To explore the local loop structure of the networks, we calculate the abundance of observed 4-node motifs. Namely, we focus on two motifs: the star motif and the 4-cycle. To quantify their abundance, we calculate their Z-score compared to their degree-preserving randomized counterparts:
@@
where $n_{\textrm{s/c}}$ is the number of occurrences of the star and 4-cycle motifs in the original networks, and the expected value $\left\langle n_{\textrm{s/c}}\right\rangle$ and standard deviation $\sigma_{\textrm{s/c}}$ is estimated by creating 200 independent randomizations [25], while keeping the degree sequence fixed. Figure 2b shows the scatter plot of $z_{\textrm{s}}$ and $z_{\textrm{c}}$ for the 15 networks. As expected, in tree networks containing no cycles the star motif is slightly over-represented $z_{\textrm{s}}>0$, and the cycle motif is slightly under-represented $z_{\textrm{s}}<0$, while for networks containing cycles, we find the opposite. The highest abundance of 4-cycles is observed for the networks representing different brain regions of the fruit fly brain.
@@
Figure 2: Abstract network properties (a) The abstract networks are composed of terminal and bifurcation points; therefore, their degree distribution is mostly concentrated on $k=1$ and $k=3$. More specifically, lattice-like networks, such as vascular and mitochondrial networks, are mostly made up of branching nodes (degree $k=3$). For the rest of the network, nodes with $k=1$ and $k=3$ are approximately evenly split, as expected for tree networks. (b) We calculate the $z$-scores of four-node star and cycle motifs of the original networks compared to random networks with the same degree distribution. Markers are the average of $n_{\textnormal{t}}=200$ independent randomizations, and error bars representing the standard error of the mean are smaller than the marker size. Both axes are symlog axes, i.e., linear from -1 to 1, log otherwise for both positive and negative values. The star motif is slightly over- and the cycle motif is slightly under-represented in tree networks (green markers), as expected. In contrast, stars are under and cycles are overrepresented in both lattices (blue markers) and linked trees (red markers). The high $z$-scores observed for the 4-cycle indicate a lattice-like structure.
@@ IV Physical properties
In the previous section, we measured properties of the abstract network capturing the connectivity of the 15 physical networks. We continue our investigation focusing on the physical properties of the system: we characterize the three-dimensional shape of the system without considering the abstract network.
@@ IV.1 Space filling and fractal dimension
A fundamental property of a physical network is its space-filling, i.e., the amount of volume it occupies from the available space. We expect that networks that are tightly packed in space are strongly affected by physicality [5, 26, 8], although the study of random physical networks built from straight links suggests that volume exclusion can play a significant role even for diminishing small space filling [26]. Many real physical networks have irregular shapes, hence much of their bounding box is unoccupied. Therefore, instead of measuring space-filling globally, we divide the axis-aligned bounding box of each physical network into rectangular boxes. We then measure the local space-filling in each box $i$ as
$\phi(i)=\frac{V_{\textrm{occ}}(i)}{V_{\textrm{box}}},$ (3)
@@
where $V_{\textrm{occ}}(i)$ is the volume of the intersection of the network and box $i$ and $V_{\textrm{box}}$ is the volume of the rectangular box. The distribution of $\phi(i)$ depends on the choice of $V_{\textrm{box}}$; therefore, to ensure consistency, we set them separately for each data set such that every bounding box is split into a $10\times 10\times 10$ grid of boxes (see SI S4.1.). Figure 3a shows the distribution of $\phi$ for all 15 physical networks, revealing that the physical networks fill out the space with mostly sparse regions and fewer denser regions. Therefore we expect that physicality will also affect the network structure unevenly: volume exclusion can limit the number and shape of links in dense regions.
@@
To further characterize the shape of the networks, we calculate their box-counting fractal dimension $D_{\textnormal{f}}$, which compactly describes the space-filling of a physical object on multiple scales [27, 28, 29] and is widely used to characterize the shape of complex biological systems [29, 30]. Possible values of $D_{\textnormal{f}}$ for connected networks range between 1 and the embedding dimension $D=3$, Fig. 3b shows $D_{\textnormal{f}}$ for the 15 networks, we find that both lattices and linked trees have $D_{\textnormal{f}}\gtrapprox 2$, while trees are typically characterized by $D_{\textnormal{f}}\lessapprox 2$, except for the anthill imprint. The fractal scaling spans at least two orders of magnitude of length scales (see SI S4.2.), again pointing towards regions of high and low physical density at different resolution levels.
@@
Figure 3: Space-filling and fractal dimension. (a) The boxplots of local space-filling $\phi$ show the coexistence of mostly sparse and fewer dense regions, where the vertical lines indicate the median, the boxes span the $25$th-$75$th percentile range, and the whiskers extend to 1.5 times the interquartile range. (b) The fractal dimension of physical networks $D_{\textrm{f}}$ ranges between $1$ and the embedding dimension $D=3$. Linked trees (red) and lattices (blue) have fractal dimension values between $D_{\textrm{f}}\approx 2.0$ and $D_{\textrm{f}}\approx 2.3$, while trees (green) have more variation, ranging from $D_{\textrm{f}}\approx 1.5$ to $D_{\textrm{f}}\approx 2.0$ (except for from the anthill imprint). We estimate $D_{\textrm{f}}$ using the box-counting method, the error bars indicate the standard deviation of the local scaling estimates (see SI S4.2).